OpenStudy (anonymous):

Can someone help me with a algebra 1 table word problem for fan and medal ?

OpenStudy (texaschic101):

put the numbers in order School A : 2,4,22,23,24,24,25,25,26 minimum value -- (smallest number) = 2 median -- (middle number or Q2) = 24 1st q(Q1) -- (4 + 22)/2 = 26/2 = 13 3rd q(Q3) -- (25 + 25) / 2 = 50/2 = 25 Interquartile range -- (Q3 - Q1) -- 25 - 13 = 12 School B : 25,27,28,28,29,29,30,32,35 minimum value -- 25 median (Q2) = 29 Q1 -- (27 + 28) / 2 = 55/2 = 27.5 Q3 -- (30 + 32) / 2 = 62/2 = 31 IQR -- 31 - 27.5 = 3.5